### v1

# EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics

Abstract
Designing emotionally intelligent conversational systems to provide comfort and advice to people experiencing distress is a compelling area of research. Previous efforts have focused on developing modular dialogue systems that treat socio-emotional strategy prediction as an auxiliary task and generate strategy-conditioned responses with customized decoders. Recently, with advancements in large language models (LLMs), end-to-end dialogue agents without explicit socio-emotional strategy prediction steps have become prevalent. However, despite their excellence in language generation, recent studies show that LLMs’ inherent preference bias towards certain socio-emotional strategies hinders the delivery of high-quality emotional support. To address this challenge, we propose decoupling strategy prediction from language generation, and introduce a novel dialogue strategy predictor, EmoDynamiX, which models the discourse dynamics between user emotions and system strategies using a heterogeneous graph. Additionally, we make use of the Emotion Recognition in Conversations (ERC) task and design a flexible mixed-emotion module to capture fine-grained emotional states of the user. Experimental results on two ESC datasets show EmoDynamiX outperforms previous state-of-the-art methods with a significant margin. Our code is available at https://github.com/cw-wan/EmoDynamiX-v2.
EmoDynamiX: Emotional Support Dialogue Strategy Prediction by Modelling MiXed Emotions and Discourse Dynamics
Chenwei Wan1        Matthieu Labeau1        Chloé Clavel2 1LTCI, Télécom Paris, Institut Polytechnique de Paris     2INRIA Paris 1{chenwei.wan, matthieu.labeau}@telecom-paris.fr 2chloe.clavel@inria.fr
1 Introduction
Figure 1: An example of an emotional support conversation between the user and the conversational agent. Our task is to predict the next socio-emotional strategy for the agent.
Providing early intervention for individuals experiencing distress from life challenges is crucial for enabling them to transition toward positive lifestyles and, consequently, fostering a more caring society. This need has inspired the NLP community to develop effective Emotional Support Conversation (ESC) systems (Liu et al., 2021). These systems aim to alleviate the distress of help-seekers and can be seen as a first step in helping them to find healthcare professionals. Recently, with the release of multi-turn and human-evaluated ESC datasets (Liu et al., 2021; Wu et al., 2022), data-driven approaches have begun to surpass rule-based methods (Van der Zwaan et al., 2012; van der Zwaan et al., 2012).
Previous work on data-driven ESC has primarily focused on modular dialogue systems as defined by Clavel et al. (2022). Modular systems feature a three-fold workflow: recognizing, planning, and generating. Examples include Tu et al. (2022), Deng et al. (2023), Liu et al. (2021) and Cheng et al. (2022). These systems select the socio-emotional strategy based on the user’s state and generate responses conditioned on the predicted strategy. In this paradigm, a strategy planner is usually optimized with a customized language decoder in either a parallel or sequential manner. However, following the advent of Large Language Models (LLMs) with stronger generative abilities, end-to-end conversational systems have started to garner more attention (Chen et al., 2023b; Zheng et al., 2023), which implies that strategy planning transitions from an explicit process to a hidden mechanism. Despite this shift, recent studies indicate that LLMs are not sophisticated socio-emotional strategy planners, as their inherent bias towards certain strategies prevents them from balancing social-oriented and task-oriented goals. For example, a study by Abulimiti et al. (2023) shows that LLMs tend to push forward the dialogue process when strategies to amend rapport should be applied, resulting in poor conversation outcomes.
Clues for mitigating the limitations of LLM-based end-to-end dialogue systems may lie in modular dialogue systems that explicitly predict strategies. Studies across various dialogue tasks demonstrate a positive correlation between high-quality socio-emotional strategy prediction and the successful fulfillment of conversational tasks (Liu et al., 2021; He et al., 2018; Joshi et al., 2021; Raut et al., 2023). Therefore, we propose treating socio-emotional strategy prediction as an independent task (Figure 1) as previously done in Vanel et al. (2023). Our aim is to develop high-quality strategy predictors that could benefit LLM-based conversation agents in the long run.
We propose EmoDynamiX, a novel socio-emotional strategy prediction framework. EmoDynamiX introduces two key innovations: fine-grained user emotion state modelling and the use of heterogeneous graphs to capture conversational dynamics. Firstly, we make use of the Emotion Recognition in Conversations (ERC) task by designing a flexible mixed-emotion module that transfers knowledge from emotion distributions predicted by a pre-trained ERC model. Additionally, we integrate user emotional states and previous agent strategies into a heterogeneous graph, effectively capturing the transitions of user emotions and their interactions with past strategies. While graph learning is widely applied in conversational recognition tasks, it remains relatively underexplored in predictive tasks. To address this, we utilize dummy nodes (Liu et al., 2022; Scarselli et al., 2008) for role-aware information aggregation, enhancing the overall performance.
Our contributions are summarized as follows: 1) We propose a socio-emotional strategy prediction framework that effectively models conversational dynamics with fine-grained emotional states. 2) We introduce role-aware dummy nodes to improve the performance of graph-learning in our predictive task. 3) Experimental results on two public ESC datasets demonstrate that EmoDynamiX outperforms all previous baselines by a significant margin, achieving over 10% improvement on ESConv (Liu et al., 2021) and over 30% improvement on the low-resource dataset AnnoMI (Wu et al., 2022).
2 Related Works
2.1 Emotional Support Conversation
Emotional Support Conversation aims to alleviate distress by expressing empathy and providing suggestions. The goal is both social-oriented and task-oriented. Liu et al. (2021) introduced a joint-learning baseline to condition response generation with predicted strategies. Cheng et al. (2023) infer the user’s persona from dialogue history to create more personalized ESC systems. Cheng et al. (2022) and Li et al. (2024) focus on identifying strategies that optimize future user feedback.
Modelling user states is a critical topic in ESC. Previous work commonly approaches this by querying commonsense knowledge graphs (Tu et al., 2022; Deng et al., 2023; Peng et al., 2022; Zhao et al., 2023; Li et al., 2024). These queries are constructed by concatenating the current utterance with specific knowledge relations, such as xReact and xEffect. The queries are then fed into COMET (Hwang et al., 2021), a generative model pre-trained on commonsense knowledge graphs, which returns the user’s reaction (xReact) or feeling (xEffect) to the situation described in the utterance. However, commonsense knowledge is too general to capture the fine-grained emotional states entailed in specific speaker turns like "Okay". In contrast, models specialized in dialogue with context-aware architectures, such as sequential or graph-based models, trained on Emotion Recognition in Conversation (ERC) datasets, can handle these nuances more effectively. Additionally, emotions are frequently mixed in real-life situations, and contradictory emotions (like Sadness and Joy) could coexist in specific contexts (Braniecka et al., 2014). Our mixed-emotion modelling approach handles this complexity better and can model a large set of subtle emotional expressions by combining primary emotions without further human annotations (as demonstrated in Section 5.4).
EmoDynamiX features two key distinctions: (1) We provide an alternative to knowledge-based user state modelling: a mixed-emotion module based on label distributions predicted by a pre-trained ERC model (2) While previous works have explored various dialogue graph structures (Li et al., 2024; Peng et al., 2022; Zhao et al., 2023), our method incorporates discourse structure, which has been proven effective in various dialogue tasks (Chen and Yang, 2021; Li et al., 2023; Zhang et al., 2023), but remains underexplored in ESC.
2.2 Graph Learning in Conversational Tasks
Graph-based approaches have proven effective in various dialogue-related tasks. In recognition tasks, such as conversational emotion recognition and dialogue act recognition, the target speaker turn aggregates information from its neighbors according to the graph structure. Studies by Ghosal et al. (2019), Ishiwatari et al. (2020), Wang et al. (2020), Fu et al. (2023), and Shen et al. (2021) design dialogue graphs based on interactions between speaker roles. Li et al. (2023) and Zhang et al. (2023) construct dialogue graphs based on discourse dependencies parsed with a pre-trained expert model, an approach also applied by Chen and Yang (2021) and Feng et al. (2021) in dialogue summarization. Yang et al. (2023) incorporate commonsense knowledge as heterogeneous nodes. Furthermore, Hu et al. (2021) and Chen et al. (2023a) model multi-modal fusion in dialogue graphs.
In predictive dialogue tasks, such as predicting the next dialogue strategy or act, predictions must rely on global information extracted from graphs. Previous works have utilized simple readout functions, such as mean/max pooling (Joshi et al., 2021) and linear layers (Raut et al., 2023). Our approach innovates by introducing dummy nodes as special placeholders for information aggregation. While dummy nodes have been previously used in other graph-learning tasks, such as graph classification and subgraph isomorphism matching (Liu et al., 2022), they have primarily served as alternatives to readout functions, which are used for obtaining embeddings of graphs or subgraphs. We are the first to employ dummy nodes in a predictive dialogue task, which is particularly useful, since it allows to clearly model role-aware interactions with previous speaker turns.
3 Problem Formulation
The task of predicting the next dialogue strategy can be written as a multi-class classification problem: assuming a dialogue comprising $T$ speaker turns, we define the dialogue history as $H^{T}=\{U^{T},A^{T},ST^{T}\}$ where $U^{T}=\{u_{t}\}_{t=1}^{T}$ is the sequence of utterances, and each $u_{t}=\{w_{n}\}_{n=1}^{N^{t}}$ is a sequence of $N^{t}$ words. $A^{T}=\{a_{t}\}_{t=1}^{T}$is the sequence of speaker roles, with $a_{t}\in\{\texttt{user},\texttt{system}\}$. $ST^{T}=\{{st}_{t}\}_{t=1}^{T}$ is the sequence of possible strategies, but they only exist for the agent. Noting $\mathcal{S}$ the set of strategies, $I_{\texttt{user}}=\left\{t,a_{t}=\texttt{user}\right\}$ the indexes of turns where the user is speaking, and denoting similarly $I_{\texttt{agent}}$, we have that $\forall t\in I_{\texttt{agent}},{st}_{t}\in\mathcal{S}$ and $\forall t\in I_{\texttt{user}},{st}_{t}=\varnothing$. Our task is, given a fixed window size of $N-1$, to predict the strategy for agent speaker turns, which we formulate as estimating the probability distribution $\mathbb{P}(st_{t}\mid H_{t-N+1}^{t-1})$ upon $\mathcal{S}$ when $t\in I_{\texttt{agent}}$. In the next section, in order to keep our notations simple, we describe our model with indexes corresponding to estimating $\mathbb{P}(st_{N}\mid H_{1}^{N-1})$.
4 Methodology
The overview of our proposed framework is illustrated in Figure 2. Our framework comprises three main components: (1) a semantic modelling module for capturing the global semantic feature in the dialogue context; (2) a heterogeneous graph learning module, designed to capture the complex interplay between user emotions and system strategies within the dialogue history; and (3) an information fusion module, which integrates the features obtained from the previous modules to produce the prediction result.
Figure 2: The overview of our proposed model which consists of a semantic modelling module, a heterogeneous graph learning module and an information fusion module.
4.1 Semantic Modelling
To effectively capture the global semantic information in the dialogue history, we adopt a common method of representing the context in a flattened sequence format:
[…]
For each speaker turn, we use its speaker role as the separating token indicating the start of the corresponding utterance. This sequence is encoded using RoBERTa (Liu et al., 2019):
[…]
We use the embedding of the [CLS] token $\boldsymbol{C}_{\texttt{[CLS]}}$ from the last hidden layer output as the global semantic representation of the dialogue history $H_{1}^{N-1}$.
4.2 Heterogeneous Graph Learning
We propose to use a heterogeneous graph (HG) to model the dynamic interactions contained in the history $H_{1}^{N-1}$. This graph is formally defined as $\mathcal{G}=\{\mathcal{V},\mathcal{B}\}$, which are respectively the set of nodes and edges. Both can be of several types: broadly, node types correspond to past strategies adopted by the conversational agent, the user’s previous emotional states, and the strategy to be predicted; while edge types model the discourse dependencies between dialogue turns and facilitate information aggregation.
Node Types: Our heterogeneous graph contains $N$ nodes $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$, each corresponding to a speaker turn, including the target turn $N$. For user turns $i\in I_{\texttt{user}}$, $v_{i}$ is an emotion node, which encapsulate the fine-grained emotion state of the user. For user turns $i\in I_{\texttt{agent}}$, $v_{i}$ is a system strategy node, that represent the specific conversational strategy implemented by the agent. Lastly, we introduce the dummy node $v_{N}$ as a placeholder for the target utterance, aggregating information from the two other node types and their interactions.
Edge Types: The edges in our heterogeneous graph fulfill dual roles: they model the discourse dependencies between dialogue turns and facilitate information aggregation towards the dummy node. Discourse dependencies correspond to edges between and within emotion and system strategy nodes: we follow Asher et al. (2016)’s definition (which include categories such as Comment and Elaboration) and pre-train a discourse parser as proposed by Chi and Rudnicky (2022) on the multi-party discourse dataset STAC (Asher et al., 2016). Details regarding the training procedure can be found in Appendix B. We note $\mathcal{R}_{\text{Discourse}}$ the set of possible dependencies; then, $\forall(i,j)$ such that $1\leq i,j\leq N-1,\langle v_{i},v_{j}\rangle\in\mathcal{R}_{\text{Discourse}}$. We give more details on these dependencies in Appendix A. The remaining edges are aggregating information from system strategy nodes, which we call self-reference: $\forall i\in I_{\texttt{agent}},\langle v_{i},v_{N}\rangle=r_{\text{self}}$; and from user emotion nodes, which we call inter-reference: $\forall i\in I_{\texttt{user}},\langle v_{i},v_{N}\rangle=r_{\text{inter}}$. We note $\mathcal{R}=\mathcal{R}_{\text{Discourse}}\cup\{r_{\text{self}},r_{\text{inter}}\}$ the set of edge types.
We will describe next how we obtain node embeddings for these three node types, and how the different edge types affect their aggregation.
4.2.1 User Emotion Node Embedding: Mixed Emotion Method
Unlike previous work that relies on commonsense knowledge, we propose leveraging a pre-trained Emotion Recognition in Conversation (ERC) model to predict emotion distributions from user utterances. We then utilize the knowledge contained in these distributions to create embeddings for fine-grained user states using a mixed-prototype approach.
Training an ERC model
Our emotion recognition model consists of a RoBERTa encoder with an MLP classifier. To incorporate the global dialogue context while classifying individual utterances, we concatenate all utterances into a single sequence. We train our model with the DailyDialog dataset (Li et al., 2017); since it does not provide annotations for speaker roles, we use the special token [SEP] as the common delimiter between all utterances:
[…]
$\begin{gathered}\boldsymbol{C}^{ERC}=\texttt{Roberta}(\texttt{[CLS]},\texttt{<ucontext>})\\ \boldsymbol{z}^{i}=\texttt{MLP}(\boldsymbol{C}^{ERC}_{\texttt{[SEP]}_{i}})\end{gathered}$ (4)
For DailyDialog, we categorize emotions into seven groups: Ekman’s six basic emotions plus Neutral, collectively referred to as $\mathcal{E}$. Additional statistical details about DailyDialog, along with our motivation for selecting it, are available in Appendix A. Detailed information regarding the implementation and training hyperparameters can be found in Appendix B.
Mixed-emotion module:
To model the user’s emotional states, we employ a trainable emotion codebook. It takes the form of a parameter matrix $\boldsymbol{E}\in\mathbb{R}^{|\mathcal{E}|\times h}$ with $h$ the embedding size for our heterogeneous graph; each of the $|\mathcal{E}|$ vector $\{\boldsymbol{E}_{k}\}_{k=1}^{|\mathcal{E}|}$ encodes a distinct emotion. For an emotion node $v_{i}$, they are combined into a node embedding $\boldsymbol{g}_{e}^{i}$ using the adjusted emotion distribution:
$\boldsymbol{g}_{e}^{i}=\boldsymbol{p}^{i}\cdot\boldsymbol{E}$ (5)
This distribution is directly obtained through the output scores of our ERC model:
$\boldsymbol{p}^{i}=\left[\frac{\exp(z^{i}_{j}/\tau)}{\sum_{k}\exp(z^{i}_{k}/\tau)}\right]_{j=1}^{|\mathcal{E}|}$ (6)
To utilize the information in the emotion label distribution more effectively, we employ a learnable temperature parameter $\tau$; details on the impact of the initialization of $\tau$ can be found in Appendix G.
Our mixed emotion approach draws inspiration from MISC, where a mixed-strategy module is proposed to condition the response generation (Tu et al., 2022), yet MISC does not study the tuning of the label distribution $\boldsymbol{p}$: for example, in an ERC dataset where the label "Neutral" is prevalent, refining the distribution to become a little "sharper" could greatly mitigate the ambiguity in the model’s predictions, especially for underrepresented classes.
4.2.2 System Strategy Node Embedding
For a speaker turn $i\in I_{\texttt{agent}}$, the dialogue strategy information is encoded as a one hot vector $\boldsymbol{s}^{i}\in\{0,1\}^{|\mathcal{S}|}$. Strategies themselves, as emotions, are represented through a parameter matrix $\boldsymbol{S}\in\mathbb{R}^{|\mathcal{S}|\times h}$. Then, we simply obtain:
$\boldsymbol{g}^{i}_{st}=\boldsymbol{s}^{i}\cdot\boldsymbol{S}$ (7)
as the embedding of the system strategy node $v_{i}$.
4.2.3 Dummy Node Embedding
Previous work relies on aggregating heterogeneous graph information using simple readout functions and linear layers, which do not consider speaker roles and lack transparency regarding the contribution of each node to the final decision. To address this, we propose using the dummy node $v_{t}$ as a placeholder for the target of prediction, which interacts with previous speaker turns in a role-aware manner. We set the embedding of the dummy node as a parameter vector $\boldsymbol{g}_{d}\in\mathbb{R}^{h}$, hence being trainable and shared among all dialog graphs.
4.2.4 Relational Graph Attention Layers
We previously defined the initial node representations for our three node types. Then, we can employ relational graph attention networks (Busbridge et al., 2019) to update these node representations. Choosing a number $K$ of attention heads, we define a relation graph attention (RGAT) layer by defining keys, queries and values parameter matrices $\boldsymbol{W}_{K}^{(r,k)},\boldsymbol{W}_{Q}^{(r,k)},\boldsymbol{W}_{V}^{(r,k)}$ for each attention head $k$ and possible type of edge $r\in\mathcal{R}$. We begin by computing the attention weights $\alpha_{i,j}^{(r_{ij},k)}$ between $v_{i}$ and $v_{j}$ under relation type $r_{ij}=\langle v_{i},v_{j}\rangle$ using:
$\begin{gathered}a_{i,j}^{(r_{ij},k)}=\sigma(\boldsymbol{W}_{Q}^{(r_{ij},k)}\boldsymbol{g}^{i}+\boldsymbol{W}_{K}^{(r_{ij},k)}\boldsymbol{g}^{j})\\ \alpha_{i,j}^{(r_{ij},k)}=\frac{\exp(a_{i,j}^{(r_{ij},k)})}{\sum_{r\in\mathcal{R}}\sum_{m\in\mathcal{N}_{r}(i)}\exp(a_{i,m}^{(r,k)})}\end{gathered}$ (8)
where $\sigma$ denotes the LeakyReLU function and $\mathcal{N}_{r}(i)$ denotes the set of the indexes of neighbouring nodes of $v_{i}$ under the edge type $r$. The result of the multi-head attention for node $v_{i}$ is then:
$\boldsymbol{h}^{i}=\|_{k=1}^{K}\sigma(\sum_{r\in\mathcal{R}}\sum_{m\in\mathcal{N}_{r}(i)}\alpha_{i,m}^{(r,k)}\boldsymbol{W}_{V}^{(r,k)}\boldsymbol{g}^{m})$ (9)
[…]
${\boldsymbol{g}}^{(1),i}=\boldsymbol{h}^{i}+\boldsymbol{g}^{i}$ (10)
In our model, we use the embedding $\boldsymbol{g}^{(L),N}$ of the dummy node after applying $L$ RGAT layers as representation for the entire heterogeneous dialogue graph.
4.3 Next Dialogue Strategy Prediction
We concatenate our global semantic embedding $\boldsymbol{C}_{\texttt{[CLS]}}$ with the heterogeneous graph embedding $\boldsymbol{g}^{(L)}_{N}$; this combined representation is fed into a simple MLP classification layer to compute a probability distribution upon $\mathcal{S}$:
$\boldsymbol{o}={softmax}(\texttt{MLP}(\boldsymbol{C}_{\texttt{[CLS]}}\|\boldsymbol{g}^{(L),N}))$ (11)
which finally gives $\mathbb{P}(st_{N}\mid H_{1}^{N-1})={o}_{st_{N}}$. We adopt the weighted cross-entropy loss as our training objective: to address class imbalance, we adjust the loss contribution from each class based on its prevalence, weighting it in proportion inverse to its frequency in the training dataset.
5 Experiments
5.1 Experimental Setups
Datasets To make our model learn strategies that are beneficial for both social and task-oriented goals, we select two ESC datasets (in English) where dialogues have been human-evaluated and filtered to ensure that the conversational outcomes are positive and the strategies applied are socially appropriate: (i) ESConv (Liu et al., 2021), an emotional support dataset annotated by trained crowd-workers. It comprises 1,300 long dialogues, averaging 29.5 speaker turns per dialogue, and features 8 dialogue strategies. To ensure fair comparisons, we adhere to the official train/dev/test split [^1] https://huggingface.co/datasets/thu-coai/esconv; (ii) AnnoMI (Wu et al., 2022), an expert-annotated counselling dataset. It includes 133 long conversations, averaging 72.9 speaker turns per dialogue, and features 9 therapist strategies. We split this dataset randomly into train/dev/test sets in an 8:1:1 ratio. Detailed dataset statistics are provided in Appendix A.
Baselines To provide extensive comparisons, we choose baselines from three criteria: (i) SOTA LLMs for dialogue tasks with few-shot prompting: ChatGPT (OpenAI, 2023) and LLaMA3-70B (Meta, 2024); (ii) models fine-tuned on both next strategy prediction and utterance generation: BlenderBot-Joint (Liu et al., 2021); (iii) models fine-tuned for emotional support dialogue strategy prediction: MISC (Tu et al., 2022) and KEMI (Deng et al., 2023). Detailed descriptions can be found in Appendix E.
Implementation Details We implemented our proposed method using PyTorch (Paszke et al., 2019), initializing with the pre-trained weights from RoBERTa and employing the tokenization tools from Huggingface Transformers (Wolf et al., 2020). For optimization, we used the AdamW optimizer (Loshchilov and Hutter, 2019). Detailed hyperparameter settings can be found in Appendix F.
Model ESConv AnnoMI
[…]
LLaMA3-70B 15.36 18.45 8.38 9.01
+ 2 shot (random) 14.68 17.74 9.45 9.59
+ 2 shot (extractive) 17.70 21.47 9.52 9.13
ChatGPT 18.14 20.27 20.31 18.12
+ 2 shot (random) 17.24 19.91 17.02 15.72
+ 2 shot (extractive) 16.55 20.01 15.29 14.22
MISC 19.89 - - -
BlenderBot-Joint 20.91 24.50 17.90 25.10
KEMI 24.66 26.80 - -
EmoDynamiX ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$ ${}^{\text{\textdagger}}$
Table 1: Experimental results on two ESC datasets. $\mathcal{M}$ stands for Macro F1 and $\mathcal{W}$ for Weighted F1 (same for Table 2). $\spadesuit$ indicates results reported in Deng et al. (2023) and all other results are produced by ourselves. † indicates statistically significant improvement ($p<0.05$) over the best baseline.
5.2 Overall Performance
As shown in Table 1, EmoDynamiX significantly outperforms all baseline models across all metrics on both datasets. Compared to the state-of-the-art joint strategy learning model KEMI on ESConv, our method delivers more diverse (Macro F1) and accurate (Weighted F1) predictions. While KEMI leverages domain-specific actual case knowledge from HEAL Welivita and Pu (2022) to better model user emotion states compared to MISC, which relies solely on commonsense knowledge from COMET, it still falls short in capturing fine-grained emotion transitions in dialogues. Our ERC-based mixed-emotion approach addresses this gap more effectively. Furthermore, by modelling the interactions between user emotion states and historical strategies using a heterogeneous graph, our model can predict strategies that are more consistent with the overall progress of the counseling process.
Comparisons with two LLM prompting baselines reveal that using LLMs alone for emotional support dialogue prediction is severely limited by their inherent biases, even with provided examples. More analysis on the impact of few-shot learning is provided in Appendix H Extensive experiments on AnnoMI further confirm that our method excels in more professional and low-resource settings. While LLM-based methods and joint learning methods exhibit a significant drop in performance (particularly LLaMA3-70B, which failed to discern these strategies and resorted to using only a few), our method maintains its advantage, showcasing its robustness and effectiveness.
Model ESConv AnnoMI
[…]
EmoDynamiX 27.40 30.81 27.48 33.09
Flattened Context 25.04↓2.36 28.68↓2.13 22.25↓5.23 27.25↓5.84
w/o Graph 25.72↓1.68 29.31↓1.50 26.95↓0.53 29.46↓3.63
w/o Mixed Emotion 25.90↓1.50 29.45↓1.36 24.71↓2.77 30.25↓2.84
w/o Discourse Parser 26.64↓0.76 30.12↓0.69 27.04↓0.44 31.59↓1.50
w/o Dummy Node 26.67↓0.73 30.16↓0.65 26.42↓1.06 29.21↓3.88
Table 2: Evaluation results of ablation study.
Model $\mathcal{R}$ $\mathcal{S}$ $\mathcal{A}$
EmoDynamiX 14.80 20.12 16.69
Flattened Context 14.04 18.90 10.65
w/o Graph 11.3 17.79 8.98
Table 3: Ablation study on three emotion-related strategies of ESConv.
Figure 3: Case study: Dialogue history and ground truth (left); visualization of the heterogeneous graph structure (middle); attention weights of the dummy node edges (right).
5.3 Ablation Study
We conducted ablation studies on the two datasets, as shown in Table 2. Our analysis is organized around several research questions:
RQ 1: Is Emotion & Strategy beneficial for strategy prediction? We compared two simplified versions based solely on the semantic modeling module with a classification head. The first version uses only flattened dialogue history with speaker roles inserted as tags (Flattened Context). This resulted in a significant drop in both Macro F1 and Weighted F1 scores compared to EmoDynamiX. We then augmented Flattened Context with user emotions and system strategies, also inserted as tags (w/o Graph). Although this augmentation led to some improvements over Flattened Context, there remains a significant performance gap compared to EmoDynamiX. This indicates that while incorporating emotions and strategies is beneficial, the effectiveness is still limited without our heterogeneous graph.
RQ 2: Does Mixed-emotion bring better results? To answer this question, we tested the effectiveness of our mixed emotion module by embedding user emotional states with one-hot vectors, omitting the mixed-emotion module (w/o Mixed Emotion). The resulting decreases in both metrics highlight the importance of leveraging emotion distributions, not just labels, for capturing fine-grained user emotion states. It is worth noting that mixed-emotion plays a more crucial role in low-resource settings, as evidenced by the more significant performance decline on AnnoMI.
RQ 3: Do we need discourse structure? We excluded the discourse parser, connecting the nodes in simple sequential order (w/o Discourse Parser). Although this led to drops in both metrics, the declines were not substantial. We hypothesize that the domain gap between the discourse parsing dataset and the ESC datasets may limit the discourse parsing module’s effectiveness.
RQ 4: Are dummy nodes better for information aggregation? We removed the dummy node and used traditional graph readout functions by concatenating mean and max node embeddings (Joshi et al., 2021). We observed performance decreases on both datasets, with a more significant decline on AnnoMI, indicating that our dummy node design is particularly beneficial in low-resource settings.
RQ 5: How does graph learning improve our model’s emotional intelligence? We compared three emotion-related dialogue strategies in ESConv: Reflection of Feelings ($\mathcal{R}$), Self-disclosure ($\mathcal{S}$), and Affirmation and Reassurance ($\mathcal{A}$). As observed in Table 3, incorporating user emotions and previous system strategies simply as tags into the flattened context (w/o Graph) is ineffective, as it drastically decreases performance across all three emotion-related strategies compared to Flattended Context. This finding highlights that modeling conversational dynamics with heterogeneous graphs can effectively mitigate preference bias and enhance our model’s emotional intelligence.
5.4 Case Study
We illustrated a case study using a conversation snippet from the ESConv test set, as shown in Figure 3. The case involves the agent deciding which strategy to apply after the user’s emotional state has transitioned positively from Frustration (as Frustration is not a category in DailyDialog, our mixed-emotion module models it as moderated sadness with a little anger) to Joy due to the agent’s previous applied strategies Providing Suggestions and Self-disclosure. The ground truth here is to apply Affirmation and Reassurance to acknowledge the user’s positive transition and encourage the user to consolidate this positive mood. Our model made the correct prediction, and the visualization offers insights into the decision-making details. By looking at the attention weights of dummy node edges, we can observe the contribution of each node. In the early stages (shallow RGAT layer), the dummy node pays more attention to previously applied socio-emotional strategies. As the RGAT layer deepens, the dummy node starts to shift its attention to user’s emotional transition, with higher weights applied to edges connecting with the emotion state nodes (especially Frustration). In short, our graph-learning module effectively captures clues from emotion/strategy dynamics for better prediction. Additionally, the attention weights provide a certain extent of transparency to the previously "black-box" decision process.
6 Error Analysis
While our method achieved SOTA results, most of the predicted strategy labels still do not match the ground truths. We examined instances where our model failed to predict the desired strategies to provide directions for future improvement.
We discovered that correctly predicting emotion-related dialogue strategies is the most difficult challenge. For our performance on ESConv (confusion matrix in Appendix I), three emotion-related strategies (Reflection of feelings, Self-disclosure, and Affirmation and Reassurance) have an average F1 score of 22.14, which is significantly lower than the overall Macro F1 score of 27.40. Among the samples where emotion-related strategies were the true label but were wrongly predicted, 63% were categorized under content-providing strategies (Question, Providing Suggestions, Information, and Restatement or Paraphrasing) and Others. This indicates that the model struggles to choose the correct timing for pushing towards task-oriented goals, failing to recognize when users need emotional soothing rather than practical suggestions or information. If emotion-related strategies are not delivered when needed, the social-oriented goals of the dialogue are underachieved, harming the overall outcome of emotionally supportive dialogue. Therefore, it is crucial for future work to deepen the understanding of users’ fine-grained and dynamic emotional states to deliver strategies at the correct timing.
7 Conclusions
In this paper, we propose EmoDynamiX, a novel socio-emotional dialogue strategy predictor that uses heterogeneous graphs to model the conversational dynamics of user emotional states and previously applied system strategies. By using pre-trained expert models, we use the discourse structure and model the emotional states of the user in a fine-grained style. Additionally, we enable speaker role-aware information aggregation with a dummy node design. Our approach has demonstrated its effectiveness by significantly improving all baselines on two public ESC datasets. We hope that our dedicated model can be integrated into future LLM-based ESC agents to mitigate the inherent limitations of LLMs in strategy planning, ultimately providing better emotional support with better transparency.
Limitations
Limitations on Ground Truths Although the ESC datasets we use are all human-evaluated, we cannot fully ensure that no strategy other than the ground truth could have worked in the same context. However, we currently lack a reliable method to accurately model user preferences and exhaustively compare all possible strategies at each step. We believe that developing such effective user simulations is a promising direction for future work, which would enable us to identify optimal strategies for each step and even allow for multi-step planning beyond just predicting the next strategy.
More on Conversational Dynamics Our method focused solely on the socio-emotional strategies of the agent and the emotional states of the user. However, integrating the agent’s emotions and the user’s dialogue acts could enhance the system further. This compound state design might necessitate a new architecture, which we plan to explore in future work.
Generalizability to Other Languages We evaluated the effectiveness of our proposed architecture using only two English datasets. It remains to be seen whether our approach can generalize to other languages or multi-language settings. It is also worth noting that since EmoDynamiX is based on expert models pre-trained on English datasets to acquire knowledge about discourse structure and emotion recognition, it may inherit cultural biases from these datasets (Gelfand et al., 2011; Hall, 1976), potentially influencing the strategy prediction.
Limitations on Expert Modules Since Emotion Recognition in Conversation (ERC) and discourse parsing are not the primary contributions of our research, we did not investigate the impact of using different model architectures or datasets. The training and integration of a cross-domain ERC module and discourse parser will be considered in future studies.
Distance to Practical Application Although our method outperformed previous baselines significantly, its performance remains unsatisfactory. This underscores the complexity of the task and indicates that additional work is required to make the socio-emotional strategy predictor a robust component in future ESC agents.
Ethics Statement
Intent of Technology We insist that conversational AI should not be developed to replace humans. Therefore, it is crucial to maintain a clear line between AI and humans (Collectif et al., 2021). Given that the training data for conversational AI, including the two datasets we selected, are primarily curated by humans, the AI (especially those trained using end-to-end methods) may exhibit human-like behaviors. For instance, in Figure 1, the system learns to utilize Self-disclosure strategy by expressing feelings of loneliness, which is not consistent with ethical recommendations. We stress that strategies leading to such human-like behaviors should be applied cautiously and potentially restricted in real-world applications to ensure safety. We believe that our next strategy prediction approach, which allows us to explicitly set desired behaviors for AI, can provide better control over conversational AI in the future.
Data Privacy All experiments were conducted using existing datasets derived from public scientific research. Any personally identifiable and sensitive information, such as user and platform identifiers, has been removed from these datasets.
Medical Disclaimer We do not provide treatment recommendations or diagnostic claims.
Transparency We detail the statistics of the datasets and the hyper-parameter settings of our method. Our analysis aligns with the experimental results.
Appendix A Datasets
ESConv (Liu et al., 2021) utilizes eight dialogue strategies: Reflection of Feelings, Self-Disclosure, Question, Affirmation and Reassurance, Providing Suggestions, Restatement or Paraphrasing, Information, and Others. The distribution of these strategies is depicted in Figure 4. ESConv collects user feedback scores (ranging from 1 to 5) after every few speaker turns to evaluate the effectiveness of emotional support. Notably, 79.9% of the scores are above 4 (Good), indicating a high overall quality of emotional support conversations in ESConv, which successfully alleviated users’ negative moods. To ensure fair comparisons with previous baselines, we did not perform filtering, though training the strategy predictors on highly rated strategies and using poorly rated ones as negative samples could be beneficial. The top-3 topics include Ongoing depression, Job crisis and Break up with parterner.
AnnoMI (Wu et al., 2022) categorizes therapist behaviors into 4 high-level types: Reflection, Question, Input, and Other. These high-level behaviors are further broken down into 9 fine-grained strategies: Simple Reflection, Complex Reflection, Open Question, Closed Question, Information, Advice, Giving Options, Negotiation/Goal-setting, and Other. Since AnnoMI is a small dataset and has very unbalanced strategy distribution, we merged Advice, Giving Options, and Negotiation/Goal-setting into a single strategy: Provide Suggestion, which is aligned with ESConv. The distribution of these strategies is illustrated in Figure 4. AnnoMI comprises 110 (82.7%) high-quality dialogues and 23 (17.3%) low-quality dialogues. To ensure our strategy predictor learns strategies that positively impact users, we excluded all low-quality conversations, retaining only the high-quality ones. The top-3 topics in AnnoMI are Reducing alcohol consumption, Smoking cessation and Weight loss.
STAC (Asher et al., 2016) is a multi-party dialogue corpus collected from an online game. It contains 1,081 dialogues, with an average of 8.5 speaker turns per dialogue. STAC includes 16 discourse dependency categories: Comment, Clarification Question, Elaboration, Acknowledgment, Continuation, Explanation, Conditional, Question-Answer Pair, Alternation, Question-Elaboration, Result, Background, Narration, Correction, Parallel, and Contrast.
DailyDialog (Li et al., 2017) is a Emotion Recognition in Conversations (ERC) dataset collected from an English learning website, its topics are closer to everyday issues and thus better suited for ESC compared to other popular counterparts collected from TV shows (Poria et al., 2019; Zahiri and Choi, 2018) or actor performances (Busso et al., 2008). DailyDialog includes 13,118 multi-turn dialogues, with an average of 7.9 speaker turns per dialogue. The emotion labels in this dataset encompass Ekman’s six basic emotions (Anger, Disgust, Fear, Joy, Sadness, Surprise) and a Neutral class.
Appendix B Implementation Details for Submodules
Discourse Parser We followed exactly the same train/dev/test split and training hyperparameters in the original paper (Chi and Rudnicky, 2022). The initial learning rate is set to 2e-5 with a linear decay to 0 for 4 epochs. The batch size is 4. The first 10% of training steps is the warmup stage. We tested the discourse parser on the test set and got a 59.0 F1 on link and relation predictions.
ERC Module For the pre-training of the ERC module, we split the DailyDialog dataset as provided in the original repository [^2] http://yanran.li/dailydialog. The learning rate was set to 2e-5, with 500 warm-up steps and a weight decay of 1e-3. The model was trained for 12,000 steps, and the best model, determined based on performance on the validation set, was used for inference on the test split. The results on the test set were as follows: an accuracy score of 82.26, a macro F1 score of 53.0, and a weighted F1 score of 83.54.
Appendix C Definitions of Strategies in ESConv
Question: asking for information related to the problem to help the seeker articulate the issues that they face.
Restatement or Paraphrasing: a simple, more concise rephrasing of the seeker’s statements that could help them see their situation more clearly.
Reflection of Feelings: describe the help-seeker’s feelings to show the understanding of the situation and empathy.
Self-disclosure: share similar experiences or emotions that the supporter has also experienced to express your empathy.
Affirmation and Reassurance: affirm the help-seeker’s ideas, motivations, and strengths to give reassurance and encouragement.
Providing Suggestions: provide suggestions about how to get over the tough and change the current situation.
Information: provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions.
Others: other support strategies that do not fall into the above categories.
Appendix D Definitions of Strategies in AnnoMI
Question open: encourage seekers to elaborate on their thoughts, feelings, and experiences, fostering self-exploration and insight. These questions cannot be answered with a simple yes or no and help build rapport and understanding.
Question closed: gather specific information, confirm details, or clarify points with concise responses. They are less exploratory but essential for obtaining precise information and ensuring clarity in the conversation.
Reflection simple: use statements that convey understanding or facilitate seeker-supporter exchanges. Simple reflection conveys understanding of what the seeker has said and adds little extra meaning.
Reflection complex: use reflective statements that show a deeper understanding of the perspective of the seeker and add substantial meaning or emphasis to what the seeker has said.
Provide suggestion: provide suggestions (Advice, Options, Goal-Setting) about how to change, but be careful to not overstep and tell them what to do.
Provide information: provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions.
Other: exchange pleasantries and use other support strategies that do not fall into the above categories.
Prompting Template Task Description
You are an intelligent emotional support assistant dedicated to helping people cope with stress and depression. To effectively comfort your users, you must select the appropriate dialogue strategy based on the context of the conversation and the user’s emotional state. Choose from the following $x$ types of strategies:
## Strategy Descriptions ## Example 1 # Dialogue context # # Output # Example 2 Dialogue Context
supporter: (strategy) …
seeker: …
Task
Now select the appropriate dialogue strategy for the next utterance according to the task description and dialogue context above (one answer only, no description), output should be in this format: (strategy).
Figure 5: Prompting template for LLMs. Examples are optional.
Appendix E Baselines
ChatGPT (OpenAI, 2023) is an advanced language model with 175 billion parameters developed by OpenAI. Using reinforcement learning from human feedback (RLHF), it generates human-like text and excels in natural language processing tasks such as conversation and content creation. The prompting is constructed with the template in Figure 5 and corresponding strategy definitions in Appendix C or Appendix D. To facilitate few-shot learning, we built a case library using the training data and extracted two examples from this library for each inference on the test data. We employed two methods for example extraction: random selection and selection based on similarity scores computed from sentence-BERT (Reimers and Gurevych, 2019) embeddings.
LLaMA3-70B (Meta, 2024) is an instruction-tuned language model with 70 billion parameters developed by Meta. Optimized for dialogue use cases, it outperforms many available open-source chat models on common industry benchmarks. The prompting templates and few-shot approach are the same as described above.
BlenderBot-Joint (Liu et al., 2021) is based on BlenderBot (Roller et al., 2021). It jointly predicts strategies and generates responses.
MISC (Tu et al., 2022) is based on BlenderBot and integrates commonsense knowledge from COMET with a mixed strategy mechanism to simultaneously predict support strategies and generate responses.
KEMI (Deng et al., 2023) is based on BlenderBot and integrates domain-specific case knowledge from HEAL with graph querying. The queries are constructed with commonsense knowledge extracted from COMET. KEMI also simultaneously predict support strategies and generate responses.
Figure 6: Analysis on the initialized value of $\tau$.
Appendix F Hyperparameter Settings
For training EmoDynamiX, we configured the batch size to 16 and set the learning rate to 4e-6, with 500 warm-up steps and a weight decay of 1e-3. Additional hyperparameters included a dimensionality of 512 for the heterogeneous graph embeddings, an initial temperature parameter $\tau$ of 0.5 for the mixed user emotion state module, and 3 layers for relational graph attention. EmoDynamiX was trained for 3000 steps on ESConv and 1200 steps on AnnoMI, separately.
For the pre-training of the ERC module, we set the learning rate to 2e-5, with 500 warm-up steps and a weight decay of 1e-3, training the model for 12,000 steps. For the pre-training of the discourse parser, we adhered to the hyperparameter settings detailed in Chi and Rudnicky (2022). All training procedures were conducted on a single Nvidia Tesla P100 GPU.
Appendix G Impact of the Initialization of $\tau$
The temperature parameter $\tau$ adjusts the shape of the probability distributions used in our mixed-emotion module. To understand the impact of the initial value of $\tau$, we performed extensive experiments, and the results are displayed in Figure 6. We observed that while both "sharpening" (lower $\tau$) and "softening" (higher $\tau$) the distribution can positively impact the overall model performance, outperforming the original distribution ($\tau=1$), "sharpening" the distribution makes the best results. This differs slightly from our expectations, as soft probabilities are typically more advantageous in learning paradigms like knowledge distillation (Hinton et al., 2015), which uses distributional knowledge to transfer learning from teacher to student models. We speculate that this outcome might be influenced by the data distribution in the DailyDialog dataset. Since "Neutral" comprises 83% of the labels, it is usually the largest or second largest category in label distributions. If the label is not "Neutral," "Neutral" plays a significant role, modulating the level of the primary emotion label, such as joy or anger. Conversely, when the label is "Neutral," the second largest category provides additional information, like "Neutral" with a hint of anger or sadness. Our results indicate that emphasizing the primary emotion category while retaining the contribution of the second-highest "moderator" category is beneficial to the overall predictive performance.
We also explored more extreme settings. When $\tau$ approaches 0, the distribution resembles a one-hot vector, leading to performance similar to the model without the mixed-emotion module (indicated by the red line in the figure). Conversely, initializing $\tau$ too high (100 or more) over-softens the distributions, allowing minor classes to introduce noise, which results in a decline in performance compared to lower $\tau$ values.
Figure 7: Impact of extractive few-shot learning on ChatGPT. Left: predictions of 0 shot; right: predictions of 2 shot.
Figure 8: Impact of extractive few-shot learning on LLaMA-70B. Left: predictions of 0 shot; right: predictions of 2 shot.
Figure 9: Normalized confusion matrix of EmoDynamiX on the test set of ESConv.
Appendix H Preference Bias of LLMs
Further analysis on ESConv indicates that in-context learning does not always positively impact large language models (LLMs) in mitigating their inherent preference bias in dialogue strategy prediction. As illustrated in Figure 7, ChatGPT (0-shot) and ChatGPT (2-shot) demonstrate different outcomes. ChatGPT effectively understands the meanings of strategies by looking at definitions alone. However, when provided with examples, whether randomly selected or extracted using sentence-BERT, its bias was actually reinforced. Conversely, the situation is different with LLaMA3-70B. As shown in Figure 8, LLaMA3-70B relied heavily on a few strategies in the 0-shot setting, indicating its limited ability to understand dialogue strategies from definitions alone. However, its strong in-context learning ability helped mitigate this issue, as it began to predict more diverse dialogue strategies when given extracted examples. It is important to note, however, that randomly selecting examples had a negative effect on LLaMA3-70B, as demonstrated in Table 1.
Appendix I Confusion Matrix of EmoDynamiX on ESConv
See Figure 9.